in 9 minutes, a conveyor belt can move a 1.00 bushels of corn into a storage bin. A smaller belt can move 1.00 bushels of corn to the storage bin in 14 minutes. If both belts are used, how long will it take to move 1,000 bushels of corn to the storage bin?
Answer by ptaylor(2198)

Let x=amount of time it takes both conveyor belts working together to move 1000 bushels of corn
Then both conveyor belts working together moves corn at the rate of 1000/x bushels per minute
Large conveyor belt moves corn at the rate of 1/9 bushel per minute
Small conveyor belt moves corn at the rate of 1/14 bushel per minute
Sooo, our equation to solve is
1/9 +1/14 =1000/x multiply both sides by 126x
14x+9x=126,000
23x=126,000
x=(126,000/23) minutes=~~~~ 5478.261 minutes
CHECK---We'll deal in fractions. That will eliminate round off errors.
In (126,000/23) minutes, large conveyor moves (126,000/23)(1/9) bushels of corn
In (126,000/23) minutes, small conveyor moves (126,000/23)(1/14) bushels of corn
(126,000/23)(1/9)+(126,000/23)(1/14)=?1000 bushels
(14000/23)+(9000/23)=1000
23000/23=1000
1000=1000
Hope this helps--ptaylor